Estimate the Debye temperature of gold if its atomic weight is 197, the density is 1.9 x 104 kg/m3 and the velocity of sound in it is 2100 m/s.
1
Expert's answer
2021-10-28T08:52:45-0400

The Debye temperature is


"T_D=\frac{hv}{\lambda},"

where


"\lambda=\sqrt[3]{\frac{m_a}{\rho}}=\sqrt[3]{\frac{197·1.66·10^{-27}}{19000}}=2.58·10^{-10}\text{ m},"

"T_D=\frac{6.63·10^{-34}·2100}{2.58·10^{-10}}=5.39·10^{-21}\text{ K}."